Find the sum of the terms between the 102nd and the 190th terms in the series

"1\/10 + 1\/5 + 3\/10 + 2\/5 + 1\/2 + ...."


1
Expert's answer
2021-02-03T01:28:55-0500
"a_1=\dfrac{1}{10}, a_2 =\dfrac{1}{5}=\dfrac{2}{10}, a_3=\dfrac{3}{10},"

"a_4=\dfrac{2}{5}=\dfrac{4}{10}, a_5 =\dfrac{1}{2}=\dfrac{5}{10}, ..."

"\dfrac{1}{10}+\dfrac{1}{5}+\dfrac{3}{10}+\dfrac{2}{5}+\dfrac{1}{2}+...=\dfrac{1}{10}\displaystyle\sum_{i=1}^\infin i"

We know that


"\displaystyle\sum_{i=1}^n i=\dfrac{n(n+1)}{2}"

Then


"\displaystyle\sum_{i=1}^{101} i=\dfrac{101(101+1)}{2}=5151"

"\displaystyle\sum_{i=1}^{190} i=\dfrac{190(190+1)}{2}=16235"

"\displaystyle\sum_{i=1}^{190} i-\displaystyle\sum_{i=1}^{101} i=16235-5151=11084"

The sum of the terms between the 102nd and the 190th terms is


"\dfrac{1}{10}(11084)=1108.4"